# UVGS: Reimagining Unstructured 3D Gaussian Splatting using UV Mapping

### v1

@@ Abstract
Figure 1: We propose UVGS - an structured image-like representation for 3DGS obtained by spherical mapping of 3DGS primitives. The obtained UVGS maps can be further squeezed to a 3-channel “3D-aware” Super UVGS image capable of bridging the gap between 3DGS and existing image foundation models. We show Super UVGS can be used to compress the 3DGS assets using pretrained image Autoencoders, and to directly generate unconditional and conditional 3DGS objects using diffusion models. To this end, we present one of the first experiments on 3DGS inpainting.
@@ 1 Introduction
The creation of high-quality 3D content is essential in applications like virtual reality, game design, robotics, and film production, where realistic 3D representations play a critical role. Typical 3D representation approaches like Neural Radiance Fields (NeRF) [28] have shown promise but require substantial computational resources, limiting their scalability for real-time applications. Moreover, NeRF is an implicit representation, which makes editing and manipulation challenging. Recently, 3D Gaussian Splatting (3DGS) [18] has emerged as a compelling alternative, enabling efficient and high-fidelity 3D rendering through a large set of Gaussian primitives that model spatial and visual properties. As an explicit representation, 3DGS offers the advantage of easier manipulation and direct access to individual primitives. However, while 3DGS offers benefits in terms of speed and visual quality, its unstructured, permutation-invariant nature presents significant challenges for generative tasks. Much like point clouds, 3DGS lacks a coherent spatial structure, impeding its integration with conventional image-based generative models. This lack of structure and coherence among 3D Gaussians hinders the application of image-based generative models [26, 65, 59], which rely on structured data representations.
Previous methods have tackled these challenges by transforming 3DGS data into structured formats, such as voxel grids [59, 9, 15] or image-based representations like Splatter Image [45] or triplanes [68]. Other approaches employ diffusion models to directly predict 3DGS attributes [31]. These methods, while achieving impressive visual results, often require substantial computational resources, memory-intensive multi-view rendering, complex architectures, which limits their scalability and flexibility for high-fidelity generation. Generating and processing unstructured 3D Gaussian directly by efficiently utilizing modern generative models like Variational Autoencoders (VAEs) and diffusion models is limited as the neural networks are not permutation invariant.
To address these shortcomings, we introduce UV Gaussian Splatting (UVGS), which provides a structured transformation of 3D Gaussian primitives into a 2D representation while preserving essential 3D information. We use spherical mapping [42] that inscribes Gaussian splats in a spherical surface, and projects attributes like position, rotation, scale, opacity, and color into an organized 14-channel image-like UV map. This mapping introduces spatial structure, resolving issues of permutation invariance by introducing local correspondences between neighboring Gaussians and global coherence across the entire 3D object. The result is a representation that functions as a “3D representation” structured in a 2D map format, enabling compatibility with powerful image-based neural network architectures.
@@
  • Efficient Structured Representation of 3DGS: We present UVGS, an image-like representation that solves permutation invariance and unstructured nature of discrete 3DGS through spherical mapping, making direct feature extraction possible by organizing unordered points into a coherent 2D representation compatible with 2D neural networks.
@@ 2 Related Work
3D Generative/Reconstruction Models for Objects: Generation or reconstruction of 3D assets has been a long standing task [34, 23, 55, 39, 27, 58, 44, 48, 4, 38, 46, 30, 29]. Previous reconstruction approaches including NeRF [28, 2] are often slow and does not provide a defining geometry [27, 39, 22, 66, 6, 35, 12, 53, 17, 25, 64]. Advancements in the field led to the emergence of explicit voxel grid based representations that encode colors and opacities directly [9, 32, 50]. These approaches achieve significant speed ups compared to the NeRF based approaches, but they couldn’t produce high fidelity assets due to the low resolution of voxel grids. On the other hand, triplane representation [14, 67, 43, 52] provides a trade-off between the quality and memory utilization. Recently, there has been a notable advancement in 3D Gaussian Splatting (3DGS) for the representation of objects and scenes that has led to the emergence of 3DGS showcasing impressive real-time results in reconstruction and generation tasks [18, 56, 21, 58, 47, 54, 33, 57, 67, 60, 15]. There has also been some recent advances in the 3D generative models for asset synthesis using the existing geometries like NeRF, voxel grids, or triplane geometries [47, 57, 67, 60, 13, 3] leveraging generative models [40, 16] and the existing 3D datasets [5, 10]. However, most of the works employing 3DGS or other representations use multiview rendering and Score Distillation Sampling (SDS) to achieve convincing generation and reconstruction capabilities [8, 34]. These approaches demand high memory and compute resources and are often quite slow in optimizing due to per scene optimization.
@@
Giving Structures to Discrete Gaussians: Although, 3DGS has led to breakthrough in the reconstruction field by demonstrating superior performance in multiple domains, the generation of 3DGS directly remains a challenge due to its discreteness and unstructured nature [65, 59]. These characteristics present substantial challenges when integrating them with conventional computer vision models, such as Autoencoders and generative models [65]. The research topic in direct learning of trained 3DGS primitives is largely unexplored [26]. Some efforts attempt to address this by directly predicting 3DGS attributes using diffusion models [31] while others like Splatter Image [45] project Gaussian objects into image-based representations through direct 3D-unaware projection. These methods struggle with maintaining multiview consistency, as the model only infers seen poses correctly, while hallucinating for unseen poses.
@@
Concurrent works [59, 15] follow the voxel-based representations to transport Gaussians into structural voxel grids with volume generation models for generating Gaussians. However, these methods are computationally expensive for high-resolution voxels, face difficulties in preserving high-quality Gaussian reconstructions due to information loss during voxelization. DiffGS [65] tries to solve the above issues by proposing three continuous functions to represent 3DGS. However, it is limited to only category-level generation and learning generic probability functions for all the categories poses significant compute and design challenges. In contrary, we introduce an efficient way to give structures to the unstructured and discrete Gaussians by taking inspiration from the developments in 3D graphics. Our method does not require any learning to map an unstructured set of Gaussians to this efficient and structured representation (termed UVGS). The proposed UVGS representation provides local and global correspondence among different Gaussian points making the widely available existing computer vision frameworks learn and extract underlying features from them.
@@ 3 Methodology
Preliminaries: 3D Gaussian Splatting (3DGS) represents a 3D object or a scene with a collection of Gaussians with different attributes to model the geometry and view-dependent appearance. For a 3DGS set, $G=\{g_{i}\}_{i=1}^{N}$, representing an object with $N$ individual Gaussians, the geometry of the $i^{th}$ Gaussian is explicitly parameterized via a 3D covariance matrix $\Sigma_{i}$ and it’s center $\sigma_{i}\in\mathbb{R}^{3}$ as: $g_{i}(x)=e^{(-\frac{1}{2}(x-\sigma_{i})^{T}\Sigma^{-1}(x-\sigma_{i}))}$ where, the covariance matrix $\Sigma_{i}=r_{i}s_{i}s_{i}^{T}r_{i}^{T}$ is factorized into a rotation matrix $r_{i}\in\mathbb{R}^{4}$ and a scale matrix $s_{i}\in\mathbb{R}^{3}$. The appearance of the $i-th$ Gaussian is represented by a color value $c_{i}\in\mathbb{R}^{3}$ and an opacity value $o_{i}\in R$. In practice, the color is represented by a series of Spherical Harmonics (SH) coefficients, but for simplicity, we represent the view-independent color by just RGB values. Thus, a single Gaussian can be represented by a set of five attributes as $g_{i}=\{\sigma_{i},~{}r_{i},~{}s_{i},~{}o_{i},~{}c_{i}\}\in\mathbb{R}^{14}$, and the entire 3DGS can be represented by a set of $N$ such Gaussians as: $G=\{\{\sigma_{i},~{}r_{i},~{}s_{i},~{}o_{i},~{}c_{i}\}\}_{i=1}^{N}$.
@@ 3.1 Spherical Mapping
Since the 3DGS is represented as a permutation invariant set with no structural correspondence among different Gaussians $g_{i}$, it becomes challenging to extract meaningful features from this set containing a few hundred thousand such points using neural networks. To address this issue, we introduce a new representation that gives structure to this unstructured set of points and solves the permutation invariance issue for faster and better feature extraction. We propose to accomplish this by employing spherical mapping to map the 3DGS object points to an image-based representation that is both invariant to random shuffling of 3DGS points and well structured. We prefer spherical mapping as others generally fails to capture the top and bottom parts of the object in the same UV map, and can introduce distortions for objects that extend far in the Z-direction.
We begin the spherical mapping by inscribing the 3DGS object into a sphere with the same center as the object in the canonical space. Inscribing a 3DGS object into a sphere involves enclosing the object within a sphere. This begins by determining the geometric center of the object. The next step is to calculate the radius of the sphere, which is achieved by measuring the Euclidean distance from the center to the farthest point on the object. The radius of the sphere is defined such that the sphere fully encloses the object. The sphere acts as a bounding volume, providing a simple geometric shape that contains the entire object.
We consider each Gaussian $g_{i}$ in 3D to be centered at the mean position represented by $\sigma_{i}$ with Cartesian coordinates $(x_{i},y_{i},z_{i})$. The aim is to get the spherical coordinates $(\rho_{i},\theta_{i},\phi_{i})$ for each Gaussian point $g_{i}$. To do so, we calculate the azimuthal $\theta_{i}$ and polar $\phi_{i}$ angles for each $g_{i}$ along with the distance from the origin to the point $\rho_{i}$. The spherical radius is defined as $\rho_{i}=\sqrt{x_{i}^{2}+y_{i}^{2}+z_{i}^{2}}$, the azimuthal angle is $\theta_{i}=\tan^{-1}(y_{i},x_{i})$, while the polar angle is $\phi_{i}=\cos^{-1}(z_{i},\rho_{i})$. The azimuthal and polar angles are then normalized, such that we can map them on a 2D UV map of $M\times N$ dimensionality with 14-channels. $\theta_{i}$ and $\phi_{i}$ are converted to degrees and mapped to UV image coordinates: $\theta_{i~{}\text{scaled}}=\left\lfloor\frac{\pi+\theta_{i}}{2\pi}\times\text{M}\right\rfloor,\;\;\phi_{i~{}\text{scaled}}=\left\lfloor\frac{\phi_{i}}{\pi}\times\text{N}\right\rfloor$ Each channel in the UV map stores the attributes of 3DGS including $\{\sigma_{i},~{}r_{i},~{}s_{i},~{}o_{i},~{}c_{i}\}\in\mathbb{R}^{14}$. We refer this 14-channel UV map as UVGS, $U\in\mathbb{R}^{M\times N\times 14}$ defined as:
@@
This transformed UVGS representation provides spatial coherence and solves the permutation invariance problem as any random arrangement of 3DGS points will now map to the same UVGS representation $U$. It should be noted that this kind of transformation will also preserve the spatial correlation between the Gaussian points in 3D and transform them to 2D UV maps by mapping them to nearby pixels. This provides both the local level correspondence among the neighboring Gaussians and the overall global correspondence for the entire object. Thus, solving the unstructured and discreteness problem. This will enable standard neural network architectures (e.g. generative models) to effectively capture correlations among neighboring Gaussians for efficient feature extraction.
@@
The Super UVGS representation effectively retains all the details of a 3DGS attributes and can be directly utilized with existing widely available image-based models. We demonstrate this by showing perfect reconstruction of 3DGS object from Super UVGS image in the experiments section. This semantically structured representation $S$ offers both local and global correspondence in representing Gaussian attributes and needs relatively less storage.
@@ 3.2 Dynamic GS Selection and Multiple Layers
When projecting 3DGS points to UV maps using spherical mapping, multiple points may map to the same pixel in UV space as shown in Fig. 3. The two 3DGS points $(g_{1})$ and $(g_{2})$ map to the same pixel on UV map $(P_{a})$ causing many-to-one mapping issue. To address this, we propose a Dynamic Selection approach where each UV pixel retains the 3DGS attributes with the highest opacity intersecting the same ray. Using the same example in Fig. 3, if opacity $o_{1}$ of Gaussian $g_{1}$ is less than opacity $o_{2}$ of $g_{2}$. then only $g_{2}$ attributes will be stored in the UV map at pixel $P_{a}$. We observed that this method helps maintain the geometry and appearance of the 3DGS object while resolving many-to-one mapping issues with minimal quality loss. For more complex objects or real-world scene representation, we stack multiple layers of UV maps, where each UVGS pixel holds the top-K opacity values of 3DGS points. This can be accomplished by inscribing the 3DGS object inside multiple spheres where each sphere maps the 3DGS attribute corresponding to the top-$K^{th}$ opacity value along the same ray. More details on this are presented in the supplementary. To show the effectiveness of proposed UVGS maps in capturing the intricacies of a complex objects, we used a pretrained image based autoencoder to reconstruct a 4 layer UVGS as shown in Fig. 4.

### v4

@@ Abstract
Figure 1: We propose UVGS - an structured image-like representation for 3DGS obtained by spherical mapping of its primitives. The obtained UVGS maps can be further unified to a 3-channel Super UVGS image to bridging the gap between 3DGS and existing image foundation models. We show Super UVGS can compress the 3DGS assets using pretrained image Autoencoders, and for direct unconditional and conditional 3DGS object generation using diffusion models. We also present one of the first experiments on 3DGS inpainting.
@@ 1 Introduction
The creation of high-quality 3D content is essential in applications like virtual reality, game design, robotics, and movie production, where realistic 3D representations play a critical role. Typical 3D representations like Neural Radiance Fields (NeRF) [nerf2021] are promising but require substantial computational resources, limiting their scalability for real-time applications. Moreover, NeRF is an implicit representation, which makes editing and manipulation challenging. Recently, 3D Gaussian Splatting (3DGS) [3dgs2023] emerged as a compelling alternative, enabling efficient and high-fidelity 3D rendering through a large set of Gaussian primitives that model spatial and visual properties. As an explicit representation, 3DGS offers several advantages over NeRF. However, while 3DGS offers benefits in terms of speed and visual quality, its unstructured, permutation-invariant nature presents significant challenges for generative tasks. Much like point clouds, it lacks a coherent spatial structure, impeding its integration with conventional image-based generative models. This lack of structure and coherence among primitives hinders the application of image-based generative models [shapesplat2024, diffgs2024, gaussiancube2024], which rely on structured data representations.
Previous methods have tackled these challenges by transforming 3DGS into structured formats, such as voxel grids [gaussiancube2024, sdfusion2023, gvgen2025] or image-based representations like Splatter Image [splatterimage2024] or triplanes [zou2024triplane]. Other approaches employ diffusion models to directly predict 3DGS attributes [gsd2024]. These methods, while achieving impressive visual results, often require substantial computational resources, memory-intensive multi-view rendering, complex architectures limiting their scalability and flexibility for high-fidelity generation. Generating and processing 3DGS directly by efficiently utilizing modern generative models like Variational Autoencoders (VAEs) and diffusion models is limited as the neural networks are not permutation invariant.
To address these shortcomings, we introduce UV Gaussian Splatting (UVGS), which provides a structured transformation of 3D Gaussian primitives into a 2D representation while preserving essential 3D information. We use spherical mapping [sphericalmapping2006] that inscribes Gaussian splats in a spherical surface, and projects attributes like position, rotation, scale, opacity, and color into an organized 14-channel image-like UV map. This mapping introduces spatial structure, resolving issues of permutation invariance by introducing local correspondences between neighboring Gaussians and global coherence across the entire 3D object. The result is a representation that functions as a “3D representation” structured in a 2D map format, enabling compatibility with powerful image-based neural network architectures.
@@
  • Efficient Structured Representation of 3DGS: We present UVGS, an image-like representation that solves permutation invariance and unstructured nature of discrete 3DGS through spherical mapping, making direct feature extraction possible by organizing unordered points into a coherent 2D representation compatible with 2D models.
@@ 2 Related Work
3D Generative/Reconstruction Models for Objects: Generation or reconstruction of 3D assets has been a long standing task [dreamfusion2022, magic3d2023, prolificdreamer2024, dreambooth3d2023, latentnerf2023, gaussiandreamer2023, dreamcraft3d2023, makeit3d2023, eg3d2022, albedogan2024, 3dfacecam2023, tsdf2024, genheld2024]. Previous reconstruction approaches like NeRF [nerf2021, mipnerf2021] are often slow and do not provide a defining geometry [latentnerf2023, dreambooth3d2023, neaf2023, udiff2024, tensorf2022, dnerf2021, plenoxels2022, neus22023, neusurf2024, geodream2023, zeroshot2024]. Advancements in the field led to the emergence of explicit voxel grid based representations that encode colors and opacities directly [sdfusion2023, diffrf2023, volumediffusion2023]. These approaches achieve significant speed ups compared to the NeRF based approaches, but they can’t produce high fidelity assets due to the low resolution of voxel grids. On the other hand, triplane representation [3dgen2023, triplanemeetsgs2024, triplanediff2023, rodin2023] provides a trade-off between the quality and memory utilization. Another line of work [geometry_img_diffusion2024, omages2024] splits the input mesh into different patches and simplifies the object generation problem to an image generation problem. However such methods rely on either cutting through the mesh to create a geometry image [geometry_img_diffusion2024] or rely on an existing UV representation of the geometry and utilize subset of existing UV islands [omages2024] resulting in loss of details. Recently, there has been a notable advancement in 3D Gaussian Splatting (3DGS) for the representation of objects and scenes leding to the emergence of 3DGS showcasing impressive real-time results in reconstruction and generation tasks [3dgs2023, 4dgs_realtimescene2024, dngaussian2024, gaussiandreamer2023, dreamgaussian2023, viewconsistentediting3dgs2025, gsedit2024, grm2024, triplanemeetsgs2024, gslrm2025, gvgen2025]. Recent advances in the 3D generative models for asset synthesis using the existing geometries like NeRF, voxel grids, or triplane geometries [dreamgaussian2023, grm2024, triplanemeetsgs2024, gslrm2025, get3d2022, difftf2023] leverage generative models [stable_diffusion, ddim] and the existing 3D datasets [shapenet2015, objaverse2023]. However, most of the works employing 3DGS or other representations use multiview rendering and Score Distillation Sampling (SDS) to achieve convincing generation and reconstruction capabilities [textto3Dusinggs2024, dreamfusion2022]. These approaches demand high memory and compute resources and are often quite slow in optimizing due to per scene optimization.
@@
Giving Structures to Discrete Gaussians: Although, 3DGS has led to breakthrough in the reconstruction field by demonstrating superior performance in multiple domains, the generation of 3DGS directly remains challenging due to its discreteness and unstructured nature [diffgs2024, gaussiancube2024]. These characteristics present substantial challenges when integrating them with conventional computer vision models, like Autoencoders and generative models [diffgs2024]. The research in direct learning of trained 3DGS primitives is largely unexplored [shapesplat2024]. Some efforts attempt to address this by directly predicting 3DGS attributes using diffusion models [gsd2024] while others like Splatter Image [splatterimage2024] project Gaussian objects into image-based representations through direct 3D-unaware projection. These methods struggle with maintaining multiview consistency, as the model only infers seen poses correctly, while hallucinating for unseen poses.
@@
Concurrent works [gaussiancube2024, gvgen2025] follow the voxel-based representations to transport Gaussians into structural voxel grids with volume generation models for generating Gaussians. However, these methods are computationally expensive for high-resolution voxels, face difficulties in preserving high-quality Gaussian reconstructions due to information loss during voxelization. DiffGS [diffgs2024] tries to solve the above issues by proposing three continuous functions to represent 3DGS. However, it is limited to only category-level generation and learning generic probability functions for all the categories poses significant compute and design challenges. In contrary, we introduce an efficient way to give structures to discrete Gaussians by taking inspiration from the developments in 3D graphics. Our method does not require any learning to map an unstructured set of Gaussians to this efficient and structured representation (termed UVGS). The proposed representation provides local and global correspondence among different Gaussian points making the widely available existing computer vision frameworks learn and extract underlying features from them.
@@ 3 Methodology
Preliminaries: 3DGS represents an object or a scene with a collection of Gaussians primitives to model the geometry and view-dependent appearance. For a 3DGS set, $G=\{g_{i}\}_{i=1}^{N}$, representing an object with $N$ individual Gaussians, the geometry of the $i^{th}$ Gaussian is explicitly parameterized via 3D covariance matrix $\Sigma_{i}$ and it’s center $\sigma_{i}\in\mathbb{R}^{3}$ as: $g_{i}(x)=e^{(-\frac{1}{2}(x-\sigma_{i})^{T}\Sigma^{-1}(x-\sigma_{i}))}$ where, the covariance matrix $\Sigma_{i}=r_{i}s_{i}s_{i}^{T}r_{i}^{T}$ is factorized into a rotation matrix $r_{i}\in\mathbb{R}^{4}$ and a scale matrix $s_{i}\in\mathbb{R}^{3}$. The appearance of the ithi-th Gaussian is represented by a color value $c_{i}\in\mathbb{R}^{3}$ and an opacity value $o_{i}\in R$. In practice, the color is represented by a series of Spherical Harmonics (SH) coefficients, but for simplicity, we represent the view-independent color by just RGB values. Thus, a single Gaussian can be represented by a set of five attributes as $g_{i}=\{\sigma_{i},~r_{i},~s_{i},~o_{i},~c_{i}\}\in\mathbb{R}^{14}$, and the entire 3DGS can be represented by a set of $N$ such Gaussians as: $G=\{\{\sigma_{i},~r_{i},~s_{i},~o_{i},~c_{i}\}\}_{i=1}^{N}$.
@@ 3.1 Spherical Mapping
3DGS is represented as a permutation invariant set with no structural correspondence among different Gaussians $g_{i}$, making it challenging to extract meaningful features from this set containing a few hundred thousands of them using neural networks. To address this, we introduce a novel representation that gives structure to this unstructured set of points and solves the permutation invariance issue for faster and better feature extraction. We propose to accomplish this by employing spherical mapping to map the 3DGS primitives to an image-like representation that is both invariant to random shuffling of 3DGS points and well structured.
We begin the mapping by inscribing the 3DGS object into a sphere with the same center as the object in the canonical space. Inscribing a 3DGS object into a sphere involves enclosing the object within a sphere. This begins by determining the geometric center of the object. The next step is to calculate the radius of the sphere, which is achieved by measuring the Euclidean distance from the center to the farthest point on the object. The radius of the sphere is defined such that the sphere fully encloses the object. The sphere acts as a bounding volume for the entire object.
We consider each Gaussian $g_{i}$ in 3D to be centered at the mean position represented by $\sigma_{i}$ with Cartesian coordinates $(x_{i},y_{i},z_{i})$. The aim is to get the spherical coordinates $(\rho_{i},\theta_{i},\phi_{i})$ for each Gaussian $g_{i}$. To do so, we calculate the azimuthal $\theta_{i}$ and polar $\phi_{i}$ angles for each $g_{i}$ along with the distance from the origin to the point, $\rho_{i}$. The spherical radius is defined as $\rho_{i}=\sqrt{x_{i}^{2}+y_{i}^{2}+z_{i}^{2}}$, the azimuthal angle as $\theta_{i}=\tan^{-1}(y_{i},x_{i})$, while the polar angle as $\phi_{i}=\cos^{-1}(z_{i},\rho_{i})$. The azimuthal and polar angles are then normalized, such that we can map them on a 2D UV map of $M\times N$ dimensionality with 14-channels. $\theta_{i}$ and $\phi_{i}$ are converted to degrees and mapped to UV image coordinates: $\theta_{i~\text{scaled}}=\left\lfloor\frac{\pi+\theta_{i}}{2\pi}\times\text{M}\right\rfloor,\;\;\phi_{i~\text{scaled}}=\left\lfloor\frac{\phi_{i}}{\pi}\times\text{N}\right\rfloor$ Each channel in the UV map stores 3DGS attributes, including $\{\sigma_{i},~r_{i},~s_{i},~o_{i},~c_{i}\}\in\mathbb{R}^{14}$. We refer this 14-channel UV map as UVGS, $U\in\mathbb{R}^{M\times N\times 14}$ defined as:
@@
This transformed UVGS representation provides spatial coherence and solves the permutation invariance problem as any random arrangement of points will now map to the same UVGS representation $U$. It should be noted that this kind of transformation will also preserve the spatial correlation between the Gaussian points in 3D and transform them to 2D UV maps by mapping them to neighboring pixels. This provides both the local level correspondence among the neighboring Gaussians and the overall global correspondence for the object. Thus, solving the unstructured and discreteness problems. This enables standard neural network architectures (e.g. CNNs) to effectively capture correlations among neighboring Gaussians for efficient feature extraction.
@@
The Super UVGS representation effectively retains all the details of 3DGS attributes and can be directly utilized with existing widely available image-based models. We demonstrate this by showing perfect reconstruction of 3DGS object from Super UVGS image in the experiments section. This semantically structured representation $S$ offers both local and global correspondence in representing Gaussian attributes and needs relatively less storage.
@@ 3.2 Dynamic GS Selection and Multiple Layers
When projecting 3DGS points to UV maps using spherical mapping, multiple points may map to the same pixel in UV space as shown in Fig. 3. Two 3DGS points $(g_{1})$ and $(g_{2})$ map to the same pixel on UV map $(P_{a})$ causing many-to-one mapping issue. To address this, we propose a Dynamic Selection approach where each UV pixel retains the 3DGS attribute with the highest opacity intersecting the same ray. Using the same example in Fig. 3, if opacity $o_{1}$ of Gaussian $g_{1}$ is less than opacity $o_{2}$ of $g_{2}$. Then only $g_{2}$ will be stored in the UV map at pixel $P_{a}$. We observed that this method helps maintain the geometry and appearance of the 3DGS object while resolving many-to-one mapping issues with minimal quality loss. For more complex objects or real-world scene representation, we stack multiple such layers of UV maps, where each UVGS pixel now holds attributes of the top-K opacity values of 3DGS primitives. This can be accomplished by inscribing the 3DGS object inside multiple spheres where each sphere maps the 3DGS attribute corresponding to the top-$K^{th}$ opacity value along the same ray. More details on this are presented in the supplementary. To show the effectiveness of proposed UVGS maps in capturing the intricacies of a complex objects, we used a pretrained image based autoencoder to reconstruct objects using a 4 layer UVGS as shown in Fig. 4.